OpenStudy (anonymous):

Two cards are drawn in succession without replacement from a standard deck of 52 cards. What is the probability that the first card is a heart given that the second card is a club? (Round your answer to three decimal places.)

OpenStudy (anonymous):

Hrart cards = 13 Club cards = 13 Total cards = 52 Probability = (13/52)(12(/51) = 3/51

OpenStudy (anonymous):

how do you turn the answer into a decimal?

OpenStudy (anonymous):

3/51=0.0588

OpenStudy (anonymous):

i did that and it said the answer was incorrect "(